# Understanding and Mitigating Miscalibration in Prompt Tuning for Vision-Language Models

### v1

@@ 6 Discussion
In this paper, we primarily focus on prompt tuning and analyze how textual divergence impacts confidence calibration. Such analysis may limit the potential scope of CLIP fine-tuning methods. To address this, we further consider a similar regularization approach based on image outliers for fine-tuning on visual representation. Specifically, we use ImageNet-1K (Deng et al., 2009) as an outlier repository and conduct experiments on four downstream datasets. For visual representation fine-tuning methods, we utilize CLIP-adapter (Gao et al., 2024) and visual prompt tuning (VPT) (Jia et al., 2022).
@@
As is shown in Table 4, we observe that visual-based DOR can successfully reduce the calibration error on new classes. For example, DOR outperforms VPT and CLIP-adapter baseline by reducing ECE by 4.64% and 1.82% on the DTD dataset, respectively. In general, visual-based DOR achieves better average calibration across various downstream datasets and leaves space for further improvement. Given that expected image outliers are not always accessible easily as text, a potential method is to generate them by diffusion (Du et al., 2024) for high-quality image outliers.
7 Related work
@@
Large pre-trained large vision-language models (Jia et al., 2021; Radford et al., 2021) have been verified that effectively comprehend visual concepts using language supervision and apply them in downstream tasks (Radford et al., 2021; Zhou et al., 2022b; Lu et al., 2022; Naeem et al., 2023; Ming & Li, 2024; Parelli et al., 2023). To further boost the downstream adaptation of pre-trained VLMs, many parameter-efficient tuning methods like vanilla prompt tuning (Zhou et al., 2022b; a; Khattak et al., 2023) and adapter tuning (Gao et al., 2024; Zhang et al., 2022) have been proposed for high efficiency. Despite the great success of CLIP fine-tuning, their effectiveness on confidence calibration has largely been overlooked, which is essential for real-world deployment.
@@
Confidence calibration has been widely studied to ensure that the confidence levels output by models accurately reflect their empirical accuracy. To achieve this, the state-of-the-art calibration methods can be categorized into regularization methods (Guo et al., 2017; Pereyra et al., 2017; Mukhoti et al., 2020) and post-hoc methods (Guo et al., 2017; Zadrozny & Elkan, 2002; 2001; Tomani et al., 2022; Xiong et al., 2023). In this work, we introduce a regularization method of fine-tuning CLIP to improve the confidence calibration, on both base and new classes.
@@
The outlier plays an important role in trustworthy machine learning research including out-of-distribution (OOD) detection (Hendrycks et al., 2019; Liu et al., 2020; Ming et al., 2022; Jiang et al., 2024; Du et al., 2024), noisy label learning (Wei et al., 2021), adversarial attack (Lee et al., 2021), long-tailed datasets re-balancing (Wei et al., 2022), etc. In this paper, we utilize text outliers to control the divergence of unseen textual distribution, and further improve the calibration of fine-tuned CLIP.
8 Conclusion & Limitation
In this paper, we introduce Dynamic Outlier Regularization (DOR), a simple yet effective technique that enhances the confidence calibration on both base and new classes. We show that current prompt tuning methods typically lead to a tradeoff between the base and new classes. Through the textual divergence, we provide a thorough explanation for the limitations of those tuning methods. By utilizing relevant but non-overlapped outlier outliers, DOR regularizes the textual distribution to preserve calibration capacity in zero-shot CLIP. Our method is compatible with existing prompt-tuning methods and can be extended to improve visual fine-tuning methods, like adapters. We hope future research can extend the insight in this work to other VLMs.
Limitations
Similar to previous regularization methods of CLIP, our method involves a hyperparameter $\lambda$ to control the weight of regularization, which will require extra computational costs for tuning. Moreover, our analysis is limited in the scope of CLIP, leaving other kinds of VLMs to be explored in the future.
Appendix A Theoretical justification
@@
To help readers understand the insights, we formalize our observations of CLIP fine-tuning that the textual divergence is a significant factor for confidence estimation in this part. As the image feature remains unchanged in the fine-tuning, the textual divergence can be translated to the variance of logits, which is computed by the similarity between the image feature and different text features. In the following, we formally show the relationship between the logit variance and the confidence.
@@
For simplicity, we consider a binary classification problem. Let $\left\{\bm{z_{i}}\right\}_{i=1}^{n}$ be a set of logit vectors (i.e., model outputs), where each vector $\bm{z}_{i}=[z_{1},z_{2}]^{\mathrm{T}}$ consists of two logits. We assume that the logit $z$ is an independent random variable drawn from a normal distribution $\mathcal{N}(\mu,\sigma^{2})$. The confidence (i.e., maximum softmax probability) $p_{i}$ is given by the softmax (or sigmoid) function defined as in Eq.(2). We have the following proposition.
Proposition 1.
@@
Appendix B Implementation Details
@@
Methods Epochs BS LR CTX
CoOp 200 32 0.002 16
CoCoOp 10 1 0.002 4
DEPT 200 32 0.002 16
KgCoOp 200 32 0.002 16
MaPLe 5 4 0.0026 2
TCP 50 32 0.002 4
CLIP-Adapter 200 32 0.002 -
VPT 5 4 0.0025 8
Our implementations are based on the open-source repository of DAC (Wang et al., 2024). Generally, We use CLIP ( ViT-B/16) (Radford et al., 2021) as the pre-trained VLM throughout our experiments and report results averaged over 3 runs. We fine-tune the model with 16 samples per class in a few-shot setting (Zhou et al., 2022a). Following the corresponding official implementation, We list the general hyperparameters in Table 6. Here, we briefly introduce the corresponding exclusive hyperparameters of each VLM tuning method. All the methods are adopted from their official implementation. For CoOp and CoCoOp, they do not contain other hyperparameters. For KgCoOp, we set $\lambda=8.0$. For MaPLe, we set prompt depth $J$ to 0 and the language and vision prompt lengths to 2. For DePT, the learning rate for updating the parameters in the devised CAT head is set to $6.5\times\delta$. where $\delta$ is the adopted learning rate of CoOp. Moreover, the weight in the linear probe is set to 0.7. For TCP, the weight for prompt fusion is 1.0, and the loss weight is the same as KgCoOp. For CLIP-adapter, we set $\alpha$ to 0.6, which is a trade-off hyperparameter between fine-tuned and zero-shot visual representation. Finally, following MaPLe, we set the context length to 8.0 and prompt depth to 12 for VPT.
Appendix C Outiler Selection
As is shown in Table 7, we can conclude that the selected outlier meets our requirement. For example, if our base class contains certain aircraft models such as ’A340-600’, ’A380’, and ’ATR-72’, the outliers we selected include words like ’air_transport’, ’air_transportation’, and ’military_plane’. These nouns are highly relevant to the downstream task but do not overlap with the base class. For further influence, we show that using outliers that are relevant but do not overlap with our fine-tuning task are helpful in reducing calibration error while preserving performance in the base classes. To verify it, we empirically demonstrate that using base classes as the regularization may sacrifice its accuracy and calibration in Figure 4. Moreover, as is shown in Table  5, dynamic text is better than fixed text since the fixed number of text may cause the model to overfit them.
@@ Appendix C Outiler Selection
Appendix D Detailed Experimental Results of fine-tuned VLMs
In this section, We present the detailed results of accuracy and Expected Calibration Error (ECE) to verify the effectiveness of our proposed DOR in Table 8(b)-9(b).
D.1 ECE
@@
D.2 Accuracy
Table 9: Accuracy (%) comparison of existing prompt tuning in the base-to-new generalization.
ZeroshotCLIP 97.16 91.33 63.57 71.79 90.06 27.73 69.36 53.01 57.00 70.99 72.39 69.49
CoCoOp 97.79 94.84 70.46 95.22 90.55 35.71 79.45 77.20 87.48 81.64 75.93 80.57
CoCoOp+DOR 97.87 95.18 69.72 93.41 90.42 34.61 79.02 75.43 86.16 81.28 75.70 79.89
CoOp 98.21 93.85 79.70 98.01 87.72 42.08 80.29 80.71 92.01 84.18 75.89 82.97
CoOp+DOR 98.13 94.63 78.11 97.44 89.78 42.26 81.30 79.98 91.39 85.54 76.61 83.20
DEPT 98.15 93.43 80.01 98.48 89.76 44.10 81.33 82.18 91.45 85.47 76.31 83.70
DEPT+DOR 98.32 94.05 79.88 98.45 90.08 44.44 81.88 82.37 90.09 85.75 76.64 83.81
KgCoOp 97.76 94.97 75.50 96.55 90.66 37.74 80.90 81.13 89.51 84.37 76.05 82.29
KgCoOp+DOR 97.78 94.97 74.68 96.04 90.63 39.56 80.14 80.83 88.95 84.02 75.81 82.13
MaPLe 97.93 95.60 72.54 96.65 90.61 36.67 80.91 79.98 91.61 83.75 76.91 82.11
MaPLe+DOR 98.00 95.04 71.60 95.44 90.64 35.59 80.92 80.17 92.31 84.16 76.74 81.87
TCP 98.17 94.59 79.91 97.91 90.68 42.20 82.65 82.72 89.98 87.19 77.49 83.95
TCP+DOR 98.10 94.59 80.23 97.98 90.62 41.50 82.62 82.48 90.65 86.69 77.37 83.89

### v4

@@ 6 Discussion
Can the criterion of DOR be extended to visual tuning? In this paper, we primarily focus on prompt tuning and analyze how textual divergence impacts confidence calibration. Such analysis may limit the potential scope of CLIP fine-tuning methods. To address this, we further consider a similar regularization approach based on image outliers for fine-tuning on visual representation. Specifically, we use ImageNet-1k (Deng et al., 2009) as an outlier repository and conduct experiments on four downstream datasets. To avoid potential semantic overlap between the outlier and fine-tuning data, we construct the outlier set by retaining images from 50% of the classes of ImageNet-1k, which ensures these classes differ as much as possible from those used during fine-tuning. For visual representation fine-tuning methods, we utilize CLIP-adapter (Gao et al., 2024) and visual prompt tuning (VPT) (Jia et al., 2022).
@@
As shown in Table 7, we observe that visual-based DOR can successfully reduce the calibration error on new classes. For example, DOR outperforms VPT and CLIP-adapter baseline by reducing ECE by 4.64% and 1.82% on the DTD dataset, respectively. We provide an additional analysis of visual distribution in Appendix K. In general, visual-based DOR achieves better average calibration across various downstream datasets and leaves space for further improvement. Given that expected image outliers are not always accessible easily as text, a potential method is to generate them by diffusion (Du et al., 2024) for high-quality image outliers.
What makes a good regularization for CLIP fine-tuning? In the experiments, we use the relevant but non-overlapped outlier with the fine-tuning task (referred to as near-OOD). Such selection raises the question: why do we prefer to choose near-OOD? To address this, we conducted an ablation on the policies of outlier selection. We considered four types of outliers: near-OOD (ours), far-OOD, random-OOD, and new classes used at test time. The far-OOD is selected by applying the opposite operation of Eq. 7. “w/o” denotes the baseline without outlier regularization. Since target new classes are unknown during fine-tuning, we view them as an oracle, which serves as a performance upper bound for base-to-new tasks. We report the average performance on the base-to-new datasets.
We present the results in Table 8. Since we primarily fine-tune the model for a specific downstream task, selecting random data or far-OOD data may not be optimal for base-to-new evaluation. Additionally, the target class is unknown during the fine-tuning phase. Therefore, we dynamically use near-OOD as regularization data, which reduces calibration errors on new classes while preserving performance on base classes. We present several word extracted by near-OOD in Appendix E. Interestingly, the random selection can serve as a strong baseline, demonstrating the robustness of our proposed regularization item. Moreover, while the oracle can achieve impressive performance on the calibration of new classes, the fixed number of outliers may cause the model to overfit them, resulting in a decline in generalization.
7 Conclusion & Limitations
In this paper, we introduce Dynamic Outlier Regularization (DOR), a simple yet effective technique that enhances the confidence calibration on both base and new classes. We show that current prompt tuning methods typically lead to a tradeoff between the base and new classes. Through the textual divergence, we provide a thorough explanation for the limitations of those tuning methods. By utilizing relevant but non-overlapped outliers, DOR regularizes the textual distribution to preserve calibration capacity in zero-shot CLIP. Our method is compatible with existing prompt-tuning methods and can be extended to improve visual fine-tuning methods, like adapters. We hope future research can extend the insight in this work to other VLMs.
Limitations. Similar to previous regularization methods of CLIP, DOR uses a hyperparameter $\lambda$ to control the weight of regularization, which will require extra computational costs. Moreover, our analysis is limited in the scope of CLIP, leaving other kinds of VLMs to be explored in the future.
Impact Statement
Foundational model plays an important role in today’s machine learning research. These models typically show remarkable zero-shot generalization capabilities and achieve better performance on downstream tasks via fine-tuning. Unfortunately, it often comes at the cost of generalization, which posing a significant challenge in real-world applications. In this paper, we investigate the confidence calibration of vision-language models (e.g., CLIP) after fine-tuning. Our goal is to mitigate overconfidence while preserving the models’ adaptability to diverse tasks. We hope our findings can extend to larger language models or more vision-language architectures, aiming to enhance their robustness after post-training in real-world scenarios.
Acknowledgment
Shuoyuan Wang and Hongxin Wei are supported by the Shenzhen Fundamental Research Program (Grant No. JCYJ20230807091809020). We gratefully acknowledge the support of the Center for Computational Science and Engineering at the Southern University of Science and Technology for our research.
Appendix A Related work
@@
Large pre-trained large vision-language models (Jia et al., 2021; Radford et al., 2021) have been verified to effectively comprehend visual concepts using language supervision and apply them in downstream tasks (e.g. image classification (Radford et al., 2021; Zhou et al., 2022b; Lu et al., 2022; Naeem et al., 2023), knowledge-augmented retrieval(Ming & Li, 2024) and visual question answering (Parelli et al., 2023)) in a zero-shot manner. Despite VLM’s effectiveness in generalizing new visual concepts, the performance of zero-shot CLIP still lags behind the fine-tuned performance on specific downstream tasks (Zhang et al., 2024a). To further boost the downstream adaptation of pre-trained VLMs, many parameter-efficient tuning methods like vanilla prompt tuning (Zhou et al., 2022b, a; Khattak et al., 2023a) and adapter tuning (Gao et al., 2024; Zhang et al., 2022) have been proposed for high efficiency. Moreover, many regularization-based tuning methods have been proposed to preserve the generalization performance on unseen classes (Yao et al., 2023, 2024; Zhu et al., 2023; Roy & Etemad, 2023; Khattak et al., 2023b). Despite the greater success of CLIP fine-tuning, the effectiveness of safety-related evaluation like confidence calibration has largely been overlooked, which is essential for real-world deployment.
@@
Confidence calibration has been widely studied to ensure that the confidence levels output by models accurately reflect their empirical accuracy. To achieve this, the state-of-the-art calibration methods can be categorized into regularization methods and post-hoc methods. For regularization methods, they either explicitly or implicitly regularize modern neural networks to have better calibration. Although regularization methods may not designed for calibration, they generally have better calibration performance including L2 regularization (Guo et al., 2017), Entropy regularization (Pereyra et al., 2017), focal loss (Mukhoti et al., 2020), etc. On the other hand, post-hoc methods fix the output probability after the training phase. For post-hoc methods, the most representative and simple method is temperature scaling (Guo et al., 2017), which learns a single scalar for rescaling the softmax logit. ATS (Joy et al., 2023) modifies the predicted confidence by per-data-point adaptive temperature. Another type of post-hoc calibration is binning-based calibration (Zadrozny & Elkan, 2001, 2002). For instance, Mix-n-Match (Zhang et al., 2020) leverages ensemble and composition techniques to achieve data efficiency and maintain accuracy in confidence estimates. Recently, several works have explored the calibration in CLIP. (Murugesan et al., 2025) investigate the calibration of CLIP under covariate shift. Given that existing post-hoc calibration on base classes can not transfer to new classes, DAC (Wang et al., 2024) fixes the logit scale of prediction via textual deviation-informed score in a post-hoc manner. Different from them, we address the calibration issue during the fine-tuning phase. In this work, we introduce a regularization based on dynamic outliers. We demonstrate that DOR boosts the calibration performance of many existing state-of-the-art prompt tuning methods of CLIP without affecting the vanilla fine-tuning objective.
@@
The outlier plays an important role in trustworthy machine learning research including out-of-distribution (OOD) detection, noisy label learning, adversarial attack, long-tailed datasets re-balancing, etc. In OOD detection, outliers are typically used to simulate the distribution of OOD data, thereby increasing the distinction between ID data and OOD data (Hendrycks et al., 2019; Liu et al., 2020; Ming et al., 2022; Jiang et al., 2024). Recently, Dream-OOD generate the expected OOD data by diffusion (Du et al., 2024). In noisy label learning, ODNL (Wei et al., 2021) leverages outliers as dynamical noisy labels to improve model robustness against noisy labels. To address the extreme class imbalance, Open-sampling (Wei et al., 2022) re-balance class priors via open-set noisy labels. OAT (Lee et al., 2021) leverages outlier data to improve model generalization in adversarial robustness, which regularizes the softmax probabilities to be a uniform distribution for outliers. In this paper, we utilize textual outliers to control the divergence of unseen textual distribution, and further improve the calibration of fine-tuned CLIP.
Appendix B Additional analysis of the motivation
B.1 Detailed Results of textual divergence
In section 3.2, we mainly derive the reason for the miscalibration issue from CoOp and KgCoOp. specifically, We use FD score to measure the diversity of textual representation and evaluate the prompt tuning methods on the value of output confidence and logit gap. To comprehensively verify our motivation, we further compare more prompt tuning methods including CoCoOp, MaPLe, PromptSRC and CoPrompt.
As is in Figure 5(a), we can observe a similar phenomenon as we discussed in section 3.2. We find that fine-tuning can significantly increase the FD score of textual features compared to the zero-shot CLIP. Notably, this observation can extend to new classes, even if they are not explicitly optimized during the fine-tuning phase. Consequently, the gap between the maximum logit and the others widens after fine-tuning in Figure 6. Therefore, the tuned CLIP with CE loss will make softmax predictions with high confidence, which can generate higher average confidence (See Figure 5(b)). Such an increase in confidence aligns with the improved accuracy on base classes. However, it is not consistent with the nearly unchanged accuracy on new classes, which leads to overconfidence.
(a) FD score
(b) Confidence
Figure 5: Comparison between zero-shot CLIP and different prompt tuning methods on UCF101 dataset. Fine-tuned CLIP tends to have higher confidence and FD score on both base and new classes.
(a) Base
(b) New
Figure 6: Comparison between the maximum logit and the average of other logits on DTD dataset. The logit gap become wider in fine-tuned CLIP.
B.2 Why using outliers for the regularization?
To further demonstrate the superiority of outliers in the regularization for CLIP fine-tuning, we provide additional analysis in this section. Specifically, we first analyze the calibration issue from the perspective of gradient conflicts (Shi et al., 2023) and present empirical evidence to understand previous regularization terms hinder the calibration of base classes. We can decouple CLIP’s optimization objective as
$\mathcal{L}_{\text{clip}}=\mathcal{L}_{\mathrm{ce}}+\lambda\cdot\mathcal{L}_{\mathrm{reg}},$
where $\mathcal{L}_{\mathrm{ce}}$ is the cross-entropy loss for classification, and $\mathcal{L}_{\mathrm{reg}}$ denotes the regularization term. Previously proposed regularization terms include $\mathcal{L}_{\mathrm{kg}}$ (KgCoOp), $\mathcal{L}_{\mathrm{distill}}$ (CoPrompt), and $\mathcal{L}_{\mathrm{scl}}$ (PromptSRC). For reasonable comparison, we set hyperparameters to be the same including optimizer, learning rate, etc. We calculate the cosine similarity of the prompt gradients between $\mathcal{L}_{\mathrm{reg}}$ and $\mathcal{L}_{\mathrm{ce}}$ to reflect the degree of gradient conflicts.
Regularization from base classes hinders original fine-tuning objective.
As shown in Figure 8, the gradient conflict distributions for KgCoOp, CoPrompt, and PromptSRC are predominantly within the range of $[-1,0]$, which indicates a conflict with the original learning objective. Considering that CoOp with vanilla $\mathcal{L}_{\mathrm{ce}}$ is an efficient calibrator for base classes, we can infer that these regularization terms may hinder the calibration performance for base classes. As an alternative, our proposed DOR leverages outliers to construct the regularization term. We observe that the gradient conflicts for DOR are primarily concentrated within the range $(-0.1,0.1]$. Compared to previous regularization terms, it shows significantly fewer conflicts in the $[-1,0]$ range. The phenomenon supports our claim that outliers can be used in regularization without interfering with the original fine-tuning objective.
Outlier-based regularization can break the calibration trade-off.
To further illustrate the actual performance of outliers in calibration, we conducted an analysis based on KgCoOp. Since KgCoOp uses a fixed number of base classes as the regularization term, we progressively replaced these texts with textual outliers at varying proportions $[0.1,0.2...1.0]$. As shown in Figure 8, as the proportion of outliers increases, the calibration of base classes improves while the performance on new classes remains largely unaffected. Such observation strongly supports our claim that outlier can effectively mitigate the miscalibration issue on new classes while maintaining the calibration performance on the base classes.
Figure 7: The distributions of gradient conflicts (in terms of $\cos\phi$). DOR shows less gradient conflict compared with recent prompt tuning methods.
Figure 8: ECE comparison as the proportion of outliers changed. Outliers can break the calibration trade-off on base and new classes.
Appendix C Theoretical justification
@@
To help readers understand the insights, we formalize our observations of CLIP fine-tuning that the textual divergence is a significant factor for confidence estimation in this part. As the image feature remains unchanged in fine-tuning, the textual divergence can be translated to the variance of logits, which is computed by the similarity between the image feature and different text features. In the following, we formally show the relationship between the logit variance and the confidence.
@@
For simplicity, we consider a binary classification problem. Let $\left\{\boldsymbol{z_{i}}\right\}_{i=1}^{n}$ be a set of logit vectors (i.e., model outputs), where each vector $\boldsymbol{z}_{i}=[z_{1},z_{2}]^{\mathrm{T}}$ consists of two logits. We assume that the logit $z$ is an independent random variable drawn from a normal distribution $\mathcal{N}(\mu,\sigma^{2})$. The confidence (i.e., maximum softmax probability) $p_{i}$ is given by the softmax (or sigmoid) function defined as in Eq.(2). We have the following proposition.
Proposition C.1.
@@
Appendix D Implementation details
@@
CoOp CoCoOp DEPT KgCoOp MaPLe TCP CLIP-Adapter VPT
Epochs 200 10 200 200 5 50 200 5
BS 32 1 32 32 4 32 32 4
LR 0.002 0.002 0.002 0.002 0.0026 0.002 0.002 0.0025
CTX 16 4 16 16 2 4 - 8
Our implementations are based on the open-source repository of DAC (Wang et al., 2024). Generally, We use CLIP ( ViT-B/16) (Radford et al., 2021) as the pre-trained VLM throughout our experiments and report results averaged over 3 runs. We fine-tune the model with 16 samples per class in a few-shot setting (Zhou et al., 2022a). Following the corresponding official implementation, We list the general hyperparameters in Table 9. For hyperparameter $\lambda$ in DOR, we set $\lambda$ = 8.0 for CoOp and 2.0 for other fine-tuning methods. We set the number of selected dynamic outlier repository to 5000. The number of outliers in each batch is the same as the base classes. Here, we adopt these VLM tuning methods from the corresponding official implementation and briefly introduce the corresponding hyperparameters of them. All the methods are adopted from their official implementation. For CoOp and CoCoOp, they do not contain other hyperparameters. For KgCoOp, we set $\lambda=8.0$. For MaPLe, we set prompt depth $J$ to 0 and the language and vision prompt lengths to 2. For DePT, the learning rate for updating the parameters in the devised CAT head is set to $6.5\times\delta$. where $\delta$ is the adopted learning rate of CoOp. Moreover, the weight in the linear probe is set to 0.7. For TCP, the weight for prompt fusion is 1.0, and the loss weight is the same as KgCoOp. For CLIP-adapter, we set $\alpha$ to 0.6, which is a trade-off hyperparameter between fine-tuned and zero-shot visual representation. Finally, following MaPLe, we set the context length to 8.0 and prompt depth to 12 for VPT.
Appendix E A close look at selected outliers
@@
As shown in Table 10, we can conclude that the selected outlier meets our requirement. For example, if our base class contains certain aircraft models such as ’A340-600’, ’A380’, and ’ATR-72’, the outliers we selected include words like ’air_transport’, ’air_transportation’, and ’military_plane’. These nouns are highly relevant to the downstream task but do not overlap with the base class. For further influence, we show that using outliers that are relevant but do not overlap with our fine-tuning task are helpful in reducing calibration error while preserving performance in the base classes. To verify it, we empirically demonstrate that using base classes as the regularization may sacrifice its accuracy and calibration in Figure 4. Moreover, As shown in Table  8, dynamic text is better than fixed text since the fixed number of text may cause the model to overfit them.
Appendix F The detailed experimental results
In this section, We present the detailed results of Expected Calibration Error (ECE) to verify the effectiveness of our proposed DOR in Table 11(b).
@@ Appendix F The detailed experimental results
Appendix G How does DOR affect the distribution of logit and probability?
To further illustrate the influence of DOR on confidence calibration, we visualize and compare the distribution of output logit and softmax confidence score for base and new classes in Figure 9. We compare zero-shot CLIP and CoOp w/ or w/o DOR on FGVCAircraft dataset. The results verify that DOR modifies the logit distribution and confidence level. We can observe that if the model tuning With CoOp+DOR, the logit distribution of base class approximate CoOp, and the new class is similar to zero-shot CLIP, respectively. A similar phenomenon is observed in the softmax probability distribution. The results meet our vision mentioned in Section 3.1, DOR can leverage the advantages of both models, which ensure the confidence calibration on both base and new classes after fine-tuning.
Appendix H The ablations of DOR
H.1 The similarity metric in outlier selection
In the outlier selection, we use cosine similarity as the distance metric for selecting textual outliers. To assess the metric sensitivity of our proposed DOR, we conduct an ablation and use three metrics including Cosine similarity, Euclidean distance (L2), and Mahalanobis distance. We report the average calibration performance on the base-to-new datasets across various prompt tuning methods, including CoOp, MaPLe, and CoPrompt.
We present the results in Table 12. We find that our DOR framework is not very sensitive to the choice of similarity metric and consistently reduces the calibration error. Surprisingly, Euclidean distance can outperform Cosine similarity across different prompt tuning methods and achieve better harmonic mean (HM). For example, using Euclidean distance with CoOp yielded an HM of 3.83%, compared to 4.58% with Cosine similarity. Despite this, we use Cosine similarity as the default distance metric throughout this work since it is widely used in the feature selection of vision-language models (Yi et al., 2024; Mayilvahanan et al., 2024; Wang et al., 2024).
Table 12: Average calibration performance (ECE%) across 11 datasets using different similarity metrics. Calibration error is given by ×$\times 10^{-2}$. “HM” denotes the harmonic mean.
CoOp MaPLe CoPrompt
Metric Base New HM Base New HM Base New HM
w/o DOR 3.07 14.58 8.83 2.75 5.46 4.11 2.56 5.96 4.26
Cosine 2.67 6.49 4.58 2.93 4.71 3.82 2.96 4.96 3.96
Euclidean 2.92 4.74 3.83 2.83 4.63 3.73 2.92 4.78 3.85
Mahalanobis 2.83 6.73 4.78 3.05 4.86 3.96 2.96 4.76 3.86
H.2 The size of selected outlier set
We evaluate how the number of selected outliers in DOR affects the calibration performance. Specifically, we present this ablation with average ECE on the base-to-new datasets with three prompt tuning methods including CoOp, MaPLe, and CoPrompt. We vary the number of outliers $k=\{10,50,100,\ldots,20000\}$.
As shown in Figure 10, increasing the number of selected outliers leads to an evident reduction in ECE on the new classes. The performance starts to reach a point of saturation with more outliers. Notably, even setting $k=10$ yields significant calibration improvements on the new classes. For the base classes, the outliers may be fixed in the batch during the fine-tuning due to the small number, leading to poor performance due to overfitting. We can observe a similar phenomenon in Table 1 (KgCoOp) or Table 8. Hence, we suggest to set a moderate number (e.g., 5000). Furthermore, when the number of outliers is sufficiently large, DOR becomes less sensitive to the exact value of numbers, demonstrating its robustness.